# Inverse Transition Learning: Learning Dynamics from Demonstrations

### v1

@@ Introduction
Definition 2.
($\epsilon$-optimality) A policy $\pi_{\epsilon}(.|s;T)$ is $\epsilon$-optimal with respect to transition dynamics $T$ if it exclusively selects actions from the $\epsilon$-ball $\epsilon(s;T)$ for all states $s$. An action $a$ is $\epsilon$-optimal with respect to dynamics $T$ if $a\in\epsilon(s;T)$.
Definition 3.
($\epsilon$-ball property) Transition dynamics $T$ satisfy the $\epsilon$-ball property for state $s$ if $\epsilon(s;T)=\epsilon(s;T^{*})$ and all invalid actions are $\epsilon$-away from all valid actions according to $Q^{*}(s,.;T)$.
We aim to learn dynamics $T$ that enforce the $\epsilon$-ball property across all states $s$, ensuring that these dynamics can discern between valid and invalid actions.
Definition 4.
(Deterministic/stochastic-policy state) A deterministic-policy state occurs when the $\epsilon$-optimal expert $\pi_{\epsilon}(.|s;T^{*})$ selects a single action $a=a^{*}$ in state $s$. A stochastic-policy state occurs when the expert selects multiple possible actions in state $s$.
Deterministic-policy states are akin to conditions well-understood by clinicians who are certain of the best treatment, whereas stochastic-policy states resemble conditions where multiple reasonable treatments are known without clear superiority.
@@
We assume we are given the MDP $\mathcal{M}$ except for the true transition dynamics $T^{*}$ (i.e., $\mathcal{M}\setminus\{T^{*}\}$). We also have access to batch data $\mathcal{D}=\{(s_{i},a_{i},s^{\prime}_{i})\}_{i=1}^{N}$, where the data which is assumed to have been generated by the behavior policy $\pi_{\epsilon}(.|s;T^{*})$, which is $\epsilon$-optimal regarding the true dynamics $T^{*}$. Note that when $\epsilon=0$, the expert $\pi_{\epsilon}(.|.;T^{*})$ is fully optimal, and hence we would only encounter deterministic-policy states. As $\epsilon$ increases, the expert begins to act more suboptimally, leading to more stochastic-policy states.
Constraints on Transition Dynamics Given $\pi_{\epsilon}(.|.;T^{*})$
In this section, we develop a set of constraints compatible with an $\epsilon$-optimal expert with respect to the true unknown dynamics $T^{*}$. These constraints are designed to impose the desired structure when estimating the true dynamics $T^{*}$ by enforcing the $\epsilon$-ball property (3) for every state $s$ (see Theorem 1). While the true transition dynamics $T^{*}$ are unknown, we assume access to $\pi_{\epsilon}(.|.;T^{*})$ for now but explain in Constraints on Transition Dynamics Given Batch Data $\mathcal{D}$ Only and Algorithm 1 how to deal with the more realistic settings of having access to trajectories only. This allows us to determine $\epsilon(s;T^{*})$ for each state $s$ by examining the actions of $\pi_{\epsilon}(.|s;T^{*})$ with non-zero probabilities. For a policy $\pi$, the constraints for each state $s$ referred in this paper as constraints($\pi$) are divided into two sets:
@@
constraint 1($\pi$): Differentiation Between Valid and Invalid Actions For every $a\in\epsilon(s;T^{*})$ and every $a^{\prime}\notin\epsilon(s;T^{*})$:
$Q^{\pi}(s,a;T)-Q^{\pi}(s,a^{\prime};T)\geq\epsilon\\ \iff\\ R(s,a)-R(s,a^{\prime})+\gamma\big{(}T(\cdot\mid s,a)-T(\cdot\mid s,a^{\prime})\big{)}^{\top}\\ \times(I-\gamma T_{\pi})^{-1}R_{\pi}\geq\epsilon$ (8)
@@
This ensures that for every state $s$, every invalid action is at least $\epsilon$-away from every valid action with respect to $Q^{\pi}(s,.;T)$.
constraint 2($\pi$): $\epsilon$-Closeness of Valid Actions For each distinct pair $(a,a^{\prime})\in\epsilon(s;T^{*})$, the absolute difference between their $Q$ values is bounded by $\epsilon$:
$\left|Q^{\pi}(s,a;T)-Q^{\pi}(s,a^{\prime};T)\right|\leq\epsilon\\ \iff\\ \left|R(s,a)-R(s,a^{\prime})+\gamma\big{(}T(\cdot\mid s,a)-T(\cdot\mid s,a^{\prime})\big{)}^{\top}\right.\\ \left.\times(I-\gamma T_{\pi})^{-1}R_{\pi}\right|\leq\epsilon$ (9)
This ensures that for each state $s$, every action within the $\epsilon$-ball under the true transition dynamics $T^{*}$ is also $\epsilon$-close under the transition dynamics $T$.
Theorem 1.
@@
Proof.
See Theorem 1 in the Appendix. ∎
In general, given any deterministic policy $\pi$, we can utilize constraints($\pi$) to recover a transition dynamics $T$ such that $\pi^{*}(T)=\pi$, and thus recover dynamics $T$ that explain the behavior induced by such policy $\pi$.
Lemma 1.
For any $\epsilon$, if $T$ satisfies constraints($\pi^{*}(T)$), then $T$ will satisfy the $\epsilon$-ball property.
Proof.
See Lemma 1 in the Appendix. ∎
Constraints on Transition Dynamics Given Batch Data $\mathcal{D}$ Only
In practical settings, direct access to the expert policy $\pi_{\epsilon}(\cdot|s;T^{*})$ for each state $s$ is typically unavailable, necessitating reliance on the batch data only $\mathcal{D}$, which might not cover all (state, action) pairs comprehensively. We construct an estimated policy, $\widehat{\pi}_{\epsilon}(\cdot|s;T^{*})$, such that for $s\in\mathcal{D}$, $\widehat{\pi}_{\epsilon}(a|s;T^{*})$ assigns a uniform probability to all actions $a$ present in $\mathcal{D}$. For state $s\notin\mathcal{D}$, we determine the optimal action $a$ using the proposed transition model $T$ (looking at $\pi^{*}(T)$), and then assign this action $a$ in the policy used to define constraints, as detailed in line 9 of Algorithm 1.
Methodology
In this section, we demonstrate how to leverage (near) optimal data with our constraints to infer both a point estimate and a posterior distribution over the true dynamics $T^{*}$, respectively referred to as Inverse Transition Learning (ITL) and Bayesian Inverse Transition Learning (BITL).
Inverse Transition Learning In Figure 4 of Appendix, we demonstrate the potential non-convexity of the feasible region arising from the constraints (Equations 8 and 9), attributed to the inverse operation on $T$. By substituting $T^{MLE}_{\widehat{\pi}_{\epsilon}(.|.;T^{*})}$ for $T$ in these constraints, we linearize them and transform the problem into a quadratic convex optimization problem, solved efficiently using CVXPY (Diamond and Boyd 2016). The optimization formulation is:
$\min_{T}\sum_{(s,a,s^{\prime})}N_{s,a,s^{\prime}}\cdot\left[T(s^{\prime}\mid s,a)-T^{\text{MLE}}(s^{\prime}\mid s,a)\right]^{2}\\ \text{subject to}\quad\forall(s,a)\in\mathcal{D},\forall a^{\prime}\notin\epsilon(s;T^{*}):\\ \left.R(s,a)-R(s,a^{\prime})+\gamma\left(T(\cdot\mid s,a)-T(\cdot\mid s,a^{\prime})\right)^{\top}\right.\\ \left.\times\left(I-\gamma T^{\text{MLE}}_{\widehat{\pi}_{\epsilon}(\cdot\mid\cdot;T^{*})}\right)^{-1}R_{\widehat{\pi}_{\epsilon}(\cdot\mid\cdot;T^{*})}\geq\epsilon,\right.\\ \left|R(s,a)-R(s,a^{\prime})+\gamma\left(T(\cdot\mid s,a)-T(\cdot\mid s,a^{\prime})\right)^{\top}\right.\\ \left.\times\left(I-\gamma T^{\text{MLE}}_{\widehat{\pi}_{\epsilon}(\cdot\mid\cdot;T^{*})}\right)^{-1}R_{\widehat{\pi}_{\epsilon}(\cdot\mid\cdot;T^{*})}\right|\leq\epsilon$ (10)
We denote the result of this optimization as $\textit{solve}_{ITL}\left(\widehat{\pi}_{\epsilon}(.|.;T^{*}),T^{MLE}\right)$, where $T^{MLE}$ is used to linearize the constraints. Linearizing the constraints with $T^{\text{MLE}}_{\widehat{\pi}_{\epsilon}(\cdot\mid\cdot;T^{*})}$ is appropriate because it reflects the state-action space where we have data, and thus where we expect $T^{MLE}$ to be accurate. This method not only accelerates training and provides guarantees on the constraints but also empirically outperforms (Herman et al. 2016), which relies on gradient descent and can be trapped in poor local optima, as shown in Table 4.
By iteratively solving the optimization problem as detailed in Algorithm 1, we obtain a point estimate $\widehat{T}^{*}$ that satisfies the $\epsilon$-ball property for each state $s$ in $\mathcal{D}$ (see Definition 3), aligning with the $\epsilon$-optimal expert behavior ($\pi_{\epsilon}(.|.;T^{*})$).
In essence, $\widehat{T}^{*}$ represents the dynamics that best fit the batch data $\mathcal{D}$ while also explaining the $\epsilon$-optimal expert behavior induced by $\pi_{\epsilon}(.|.;T^{*})$.
@@
1:  $i\leftarrow 0$
2:  $\pi^{(0)}(\cdot|s)\leftarrow\widehat{\pi}_{\epsilon}(\cdot|s;T^{*})$ for $s\in\mathcal{D}$
3:  $\pi^{(0)}(\cdot|s)\leftarrow\text{uniform distribution}$ for $s\notin\mathcal{D}$
4:  $T^{(0)}\leftarrow T^{\text{MLE}}$
5:  $T^{(1)}\leftarrow\textit{solve}_{ITL}(\pi^{(0)},T^{(0)})$
6:  $i\leftarrow i+1$
@@
Bayesian Inverse Transition Learning Based on our constraints, we have demonstrated how to estimate the dynamics such that it aligns with an expert’s behavior. However, both constraint 1 and constraint 2 (Inequality 8 and 9) lead to an underdetermined problem with infinitely many solutions. Instead of introducing a loss function as in optimization problem (10) and baseline MCE (Herman et al. 2016), we infer a posterior distribution on the true transition dynamics $T^{*}$ to quantify such uncertainty. This approach also lets us apply the actual constraints to each sample directly, instead of approximating them through linearization as is done in point estimate setting. We introduce a sampling-based technique to infer this distribution, denoted as $P_{\epsilon}(T^{*}|\mathcal{D})$, assuming data is generated by an $\epsilon$-optimal expert. We use our constraints to ensure each sample satisfies the $\epsilon$-ball property for each state $s\in\mathcal{D}$.
HMC with Reflection for ITL To enhance sampling efficiency within our constrained, high-dimensional space, we employ Hamiltonian Monte Carlo (HMC) with reflection ((Betancourt 2011)) (See 2). A more detailed version of the algorithm along with more details about inference can be find in the Algorithm 3 of the Appendix . This method involves transformations necessary for ensuring each sample lies within the Dirichlet simplex. If samples fail to meet constraints 1 & 2 (8, 9) or exit the feasible region after the leap-frog integration, they are rejected. We start HMC with the point estimate $\widehat{T}^{*}$ that satisfies the $\epsilon$-ball property (3) for observed states. This ensures necessary initial feasibility, leading to a rejection rate of about 20 to 60% using an adaptive step size, a marked improvement over nearly 100% rejection in rejection sampling.
Algorithm 2 HMC with reflection for Bayesian ITL
1:  First momentum half step
2:  $m\leftarrow m-\frac{1}{2}\alpha\nabla E(w)$
3:  for $l=0$ to $L$ do
4:     Full spatial step
5:     $w\leftarrow w+\alpha m$
6:     Transform $w$ into $T$
7:     $T\leftarrow\text{w\_to\_T}(w)$
8:     Create $\widehat{\pi}_{\epsilon}(\cdot\mid\cdot;T^{*})$ with proposed $T$
9:     Squeeze maximum info
10:     $\pi\leftarrow\pi^{*}(T^{(i)})$
12:     if all constraints($\pi$) satisfied then
13:        Full momentum step
14:        $m\leftarrow m-\alpha\nabla E(w)$
15:     else
16:        Bounce
17:        constraints($\pi$) $c$ is violated
18:        $\hat{n}\leftarrow\nabla c(T)/\|\nabla c(T)\|$
19:        $m\leftarrow m-2(m\cdot\hat{n})\hat{n}$
20:     end if
21:  end for
22:  Full spatial step
23:  $w\leftarrow w+\alpha m$
24:  Last momentum half step
25:  $m\leftarrow m-\frac{1}{2}\alpha\nabla E(w)$
Experimental Setup

### v2

@@ 4 ESTIMATING TRANSITION DYNAMICS WITH $\epsilon$-OPTIMAL EXPERT
For each state $s$, $\epsilon(s;T^{*})$ contains the actions that are near-optimal with respect to the true dynamics $T^{*}$.
Definition 2 ($\epsilon$-optimality).
A policy $\pi_{\epsilon}(.|s;T^{*})$ is $\epsilon$-optimal with respect to the true unknown transition dynamics $T^{*}$ if it exclusively selects actions from the $\epsilon$-ball $\epsilon(s;T^{*})$ for all states $s$:
$\pi_{\epsilon}(a|s;T^{*})>0\iff a\in\epsilon(s;T^{*}),\quad\forall s\in\mathcal{S}$
Such a policy $\pi_{\epsilon}(.|s;T^{*})$ can make a mistake of at most $\epsilon$ at each time step relative to $Q^{*}(.,.;T^{*})$. We assume our demonstrations are generated by such policy.
Definition 3 ($\epsilon$-ball property).
Dynamics $T$ satisfy the $\epsilon$-ball property for state $s$ if $\epsilon(s;T)=\epsilon(s;T^{*})$ and every action $a^{\prime}\notin\epsilon(s;T)$ is at least $\epsilon$-away from all actions $a\in\epsilon(s;T)$ with respect to $Q^{*}(s,.;T)$:
$\epsilon(s;T)=\epsilon(s;T^{*})\quad\text{and}$ (6)
$Q^{*}(s,a;T)-Q^{*}(s,a^{\prime};T)\geq\epsilon,$
$\qquad\forall a\in\epsilon(s;T),\,\forall a^{\prime}\notin\epsilon(s;T)$
We assume the true dynamics $T^{*}$ satisfy the $\epsilon$-ball property for each state $s$, splitting actions into two groups: near-optimal and suboptimal. Actions inside $\epsilon(s;T^{*})$ are valid, and those outside $\epsilon(s;T^{*})$ invalid. Our goal is to learn dynamics $T$ that satisfy this property for every state $s$. Such dynamics can distinguish near-optimal actions from suboptimal ones.
Definition 4 (Deterministic/stochastic-policy state).
A deterministic-policy state occurs when the $\epsilon$-optimal expert $\pi_{\epsilon}(.|s;T^{*})$ selects a single action in state $s$ (meaning the expert is deterministically optimal). A stochastic-policy state occurs when the expert selects multiple actions in such state:
$\text{Deterministic-policy state:}\quad|\epsilon(s;T^{*})|=1$
$\text{Stochastic-policy state:}\quad|\epsilon(s;T^{*})|>1$
Deterministic-policy states are akin to conditions well-understood by clinicians who are certain of the best treatment, whereas stochastic-policy states resemble conditions where multiple reasonable treatments exist without clear superiority.
@@
We assume we are given the MDP $\mathcal{M}$ except for the true transition dynamics $T^{*}$ (i.e., $\mathcal{M}\setminus\{T^{*}\}$). We also have access to batch data $\mathcal{D}=\{(s_{i},a_{i},s^{\prime}_{i})\}_{i=1}^{N}$, where the data is assumed to have been generated by the behavior policy $\pi_{\epsilon}(.|s;T^{*})$, which is $\epsilon$-optimal with respect to the true dynamics $T^{*}$. Note that when $\epsilon=0$, the expert $\pi_{\epsilon}(.|.;T^{*})$ is fully optimal, and hence we would only encounter deterministic-policy states.
Why Access to the Reward Function is Reasonable
Unlike transition dynamics, which describe the complex physiological response of the human body to treatments and are difficult to infer even with domain knowledge, the reward function in healthcare applications is often more accessible. In the case of hypotension management, clinical expertise can reasonably define rewards that distinguish good and bad outcomes, such as stabilizing blood pressure or preventing adverse events. Therefore, in our setup, it is natural to assume access to the reward function while focusing on learning the transition dynamics $T^{*}$. However, we recognize that in other domains, such as robotics, specifying a reward function can be challenging. For such settings, we provide an extension of our algorithm in Appendix B that jointly learns the transition dynamics and the reward function simultaneously. Our results demonstrate that this joint learning approach achieves performance comparable to when the true reward function is provided.
4.1 Constraints on Transition Dynamics Given $\pi_{\epsilon}(.|.;T^{*})$
In this section, we develop a set of constraints for learning transition dynamics $T$ motivated by the $\epsilon$-optimal assumption we have on the expert (See Definition 2). We denote these constraints on the dynamics $T$ collectively as Constraints($T$), which consist of two types:
Constraint 1($T$): Separation Between Valid and Invalid Actions
@@
For every valid action $a\in\epsilon(s;T^{*})$ and every invalid action $a^{\prime}\notin\epsilon(s;T^{*})$, $T$ must satisfy:
$Q^{*}(s,a;T)-Q^{*}(s,a^{\prime};T)\geq\epsilon$ (7)
Expanding the Q-values in terms of $T$:
$R(s,a)-R(s,a^{\prime})+\gamma\big(T(\cdot\mid s,a)$
$-T(\cdot\mid s,a^{\prime})\big)^{\top}(I-\gamma T_{\pi^{*}(T)})^{-1}R_{\pi^{*}(T)}\geq\epsilon$ (8)
@@
where $\pi^{*}(T)$ denotes the optimal policy induced by dynamics $T$. This constraint ensures that for every state $s$, all invalid actions are at least $\epsilon$-away from all valid actions with respect to $Q^{*}(s,\cdot;T)$.
Constraint 2($T$): $\epsilon$-Closeness Among Valid Actions
For each pair of valid actions $(a,a^{\prime})\in\epsilon(s;T^{*})\times\epsilon(s;T^{*})$ where $a\neq a^{\prime}$, $T$ must satisfy:
$\left|Q^{*}(s,a;T)-Q^{*}(s,a^{\prime};T)\right|\leq\epsilon$ (9)
Expanding the Q-values in terms of $T$:
$\left|R(s,a)-R(s,a^{\prime})+\gamma\big(T(\cdot\mid s,a)-T(\cdot\mid s,a^{\prime})\big)^{\top}\right$.
$\quad\left.(I-\gamma T_{\pi^{*}(T)})^{-1}R_{\pi^{*}(T)}\right|\leq\epsilon$ (10)
This constraint ensures that for each state $s$, valid actions remain $\epsilon$-close with respect to $Q^{*}(s,\cdot;T)$.
Enforcing the $\epsilon$-ball property: Together, these constraints enforce the $\epsilon$-ball property (Definition 3) for the learned dynamics $T$. Constraint 2($T$) ensures that all valid actions remain $\epsilon$-close to each other, while Constraint 1($T$) ensures that all invalid actions stay at least $\epsilon$-away from valid actions with respect to dynamics $T$. This is exactly what the $\epsilon$-ball property requires. Hence, if dynamics $T$ satisfy both sets of constraints, then it satisfies the $\epsilon$-ball property.
4.2 Constraints on Transition Dynamics in practice
While the true dynamics $T^{*}$ and sets $\epsilon(s;T^{*})$ are unknown, we approximate them by collecting all actions $a$ taken in state $s$ from the dataset $\mathcal{D}$, denoted $\hat{\epsilon}(s;T^{*})$. This relies on the assumption that $\mathcal{D}$ was generated by an $\epsilon$-optimal expert under the true dynamics $T^{*}$.
In practice, we do not have direct access to the expert policy $\pi_{\epsilon}(\cdot\mid s;T^{*})$. Instead, we rely on the dataset $\mathcal{D}$, which may not cover all (state, action) pairs. We therefore define an estimated policy $\hat{\pi}_{\epsilon}(\cdot\mid s;T^{*})$: for states $s\in\mathcal{D}$ it assigns equal probability to all observed actions, while for $s\notin\mathcal{D}$ it selects the best action under the proposed transition model $T$ (using $\pi^{*}(T)$). This policy is then used to define the constraints in Algorithm 1 &\&  2, which integrates these components to estimate the true but unknown dynamics $T^{*}$.
5 METHODOLOGY
In this section, we describe how, given (near-)optimal data, our constraints can be applied to infer both a posterior distribution over the true dynamics $T^{*}$ and a point estimate. These are referred to as Bayesian Inverse Transition Learning (BITL) and Inverse Transition Learning (ITL), respectively.
Bayesian Inverse Transition Learning Our constraints (Inequalities 7 and 9) lead to an underdetermined problem with infinitely many solutions. Instead of introducing a loss function as the baseline MCE (Herman et al.,, 2016), we infer a posterior distribution on the true transition dynamics $T^{*}$ to quantify such uncertainty. We introduce a sampling-based technique to infer this distribution, denoted as $P_{\epsilon}(T^{*}|\mathcal{D})$, assuming data is generated by an $\epsilon$-optimal expert. We use our constraints to ensure each sample satisfies the $\epsilon$-ball property (See Definition 3) for each state.
HMC with Reflection To enhance sampling efficiency within our constrained, high-dimensional space, we employ Hamiltonian Monte Carlo (HMC) with reflection (Betancourt,, 2011). The log-likelihood function of our target distribution takes the following form: $\log P(T^{*}\mid\mathcal{D})=\sum_{s,a}\log\text{Dir}(\mathbf{N}_{s,a}+\delta\mid s,a)$. Applying standard HMC to Dirichlet distributions is challenging due to their simplex constraints. To address this, we use the invertible transformations from (Betancourt,, 2012), combined with a logit transformation, to map the constrained transition dynamics $T$ into an unconstrained position variable $w$.
Our sampling procedure (Algorithm 1) utilizes first-order leapfrog integration. Let $m$ denote the momentum, $\alpha$ the step size, $\nabla E(w)$ the energy gradient, and $L$ the number of integration steps. During each spatial step, we transform $w$ back to $T^{(i)}$ to compute the optimal policy $\pi^{*}(T^{(i)})$ and check Constraints 1 & 2 (Inequalities 7 & 9). If a constraint $c_{s,a,a^{\prime}}(T^{(i)})$ is violated, the trajectory “bounces” by reflecting the momentum $m$ off the constraint’s normal vector $\hat{n}$.
Algorithm 1 HMC with reflection for Bayesian ITL
@@ 5 METHODOLOGY
2:for $i=0$ to $L$ do
3:  $w\leftarrow w+\alpha m$ // Full spatial step
4:  $T^{(i)}\leftarrow\text{w\_to\_T}(w)$
5:  Compute $\pi^{*}(T^{(i)})$
[…]
7:   $m\leftarrow m-\alpha\nabla E(w)$ // Full momentum step
8:  else
9:   $\hat{n}\leftarrow\nabla c_{s,a,a^{\prime}}(T^{(i)})/\|\nabla c_{s,a,a^{\prime}}(T^{(i)})\|$
10:   $m\leftarrow m-2(m\cdot\hat{n})\hat{n}$
11:  end if
12:end for
[…]
A final clean-up step ensures that if the leapfrog integration terminates outside the feasible region, the sample is immediately rejected. This guarantees that every accepted sample satisfies the $\epsilon$-ball property for each state $s\in\mathcal{D}$. To ensure initial feasibility, we warm-start the HMC chain with the point estimate $\widehat{T}^{*}$ obtained via ITL in Algorithm 2 (See below). By reflecting off boundaries rather than blindly rejecting, our algorithm achieves a rejection rate of about 20% to 60% across our experiments—a significant improvement over the nearly 100% rejection rate of standard rejection sampling.
@@
[…]
@@
Inverse Transition Learning In addition to our posterior estimate, we can also infer a point estimate $\widehat{T}^{*}$. By introducing a quadratic loss that enforces such an estimate to remain close to $T^{\text{MLE}}$, and by linearizing Constraints 1 & 2 using a fixed reference policy $\tilde{\pi}$ and fixed transition dynamics $\tilde{T}$, we form a quadratic convex optimization problem that can be solved efficiently using CVXPY (Diamond and Boyd,, 2016):
$\min_{T}\sum_{(s,a,s^{\prime})}N_{s,a,s^{\prime}}\cdot\left[T(s^{\prime}\mid s,a)-T^{\text{MLE}}(s^{\prime}\mid s,a)\right]^{2}$
$\text{subject to}\quad\forall(s,a)\in\mathcal{D},\forall a^{\prime}\notin\hat{\epsilon}(s;T^{*}):$
$\quad R(s,a)-R(s,a^{\prime})+\gamma\left(T(\cdot\mid s,a)-T(\cdot\mid s,a^{\prime})\right)^{\top}$
$\quad\times\left(I-\gamma\tilde{T}_{\tilde{\pi}}\right)^{-1}R_{\tilde{\pi}}\geq\epsilon,$
$\text{and subject to}\quad\forall(s,a)\in\mathcal{D},\forall a^{\prime}\in\hat{\epsilon}(s;T^{*}):$
$\quad\left|R(s,a)-R(s,a^{\prime})+\gamma\left(T(\cdot\mid s,a)-T(\cdot\mid s,a^{\prime})\right)^{\top}\right$.
$\quad\left.\times\left(I-\gamma\tilde{T}_{\tilde{\pi}}\right)^{-1}R_{\tilde{\pi}}\right|\leq\epsilon$ (11)
[…]
This yields the iterative procedure described in Algorithm 2. We initialize our constraints using $\widehat{\pi}_{\epsilon}$ and $T^{\text{MLE}}$. In subsequent steps, $\textit{solve}_{ITL}(\{\pi^{(k)},T^{(k)}\}_{k=0}^{i})$ denotes solving the optimization problem with additional constraints accumulated from all iterations up to $i$. We keep augmenting the constraint set until we reach a point estimate $\widehat{T}^{*}$ that satisfies the $\epsilon$-ball property for each state $s\in\mathcal{D}$ (Definition 3). In essence, $\widehat{T}^{*}$ represents the dynamics that best fit the batch data $\mathcal{D}$ while also explaining the $\epsilon$-optimal expert behavior.
6 EXPERIMENTAL SETUP